# Stretching Beyond the Obvious: A Gradient-Free Framework to Unveil the Hidden Landscape of Visual Invariance

### v2

@@ 3.1 The Stretch-and-Squeeze algorithm (SnS)
For the characterization of invariance, we set this representation space at three distinct hierarchical levels within ResNet50: (i) $\kappa=0$, i.e., the input pixel space (denoted low_level), (ii) a mid-level convolutional layer (1st convolutional stage in layer 3, mid_level), and (iii) a deep convolutional layer (7th convolutional stage of layer 4, high_level). Our primary target units $u_{\mathrm{targ}}^{\ell}$ were chosen in the final readout layer (fully connected), as these units are selective for specific object categories, allowing us to study the invariance of the whole network to transformations of the image classes those neurons are tuned to. To demonstrate SnS’s generalizability beyond category-selective units, we also performed optimizations targeting units within mid_level and high_level, which are potentially more analogous to biological neurons in the visual system. In these cases, the input-side distance was computed in pixel space (i.e. $k=0$). For adversarial attack generation, we restricted our analysis to the configuration targeting readout layer units and defining input distance in pixel space. SnS optimizations were conducted on the units of two specific instances of ResNet50, our subject network: the standard ImageNet-pretrained model available in PyTorch (Paszke et al., 2019), and a robust counterpart generated via adversarial training with an $L_{2}$ perturbation norm constraint of $\epsilon=3$ (Engstrom et al., 2019a). SnS invariances were studied also in ResNet18 and VGG16_bn (see Section 3.2) and Vision Transformers (ViT, see Supplementary Material, Section G).
@@ 4.2 SnS reveals layer-specific invariant manifolds
Yet another way to characterize the invariance manifolds is to estimate their intrinsic dimension (ID). We did so for an example readout unit of $L_{2}$-robust ResNet50, using both PCA and 2NN-ID, a state-of-the-art nonlinear estimator (Facco et al., 2017) (see Section H in the Supplementary Material). 2NN-ID yielded ID estimates one order of magnitude lower than PCA, indicating that the geometry of the three manifolds was highly nonlinear (Fig. S10). Both methods, however, revealed a consistent ID ranking across stretching layers: lowest for the pixel space, highest for mid layers, and lower again for deep layers - a pattern mirroring previously reported trends in the ID of image representations across deep CNNs and visual cortex (Ansuini et al., 2019; Muratore et al., 2022a)
@@
Hierarchical invariances are robust to representation subsampling. To assess SnS’s applicability to neuroscience, where only a small neuronal subpopulation is typically recorded from a visual area, we generated the invariant images using heavily subsampled hidden layer representations. Both qualitative inspection and representational distance analysis revealed that the resulting invariances closely resembled those obtained using the full layer representation, supporting SnS’s potential for neuroscience experiments (see Supplementary Material, Section F for further details).
@@ 4.3 Invariant images for robust and standard networks are perceptually different
We next evaluated how recognizable the invariant images generated by SnS were by humans and other CNNs. In our tests, invariant images for both kinds of networks (i.e., standard and $L_{2}$-robust) were generated by stretching the low- (pixel), mid- and high-level representations of the same set of natural images sampled from 12 different Imagenet categories (Fig. 4a). Human subjects performed a 12-AFC classification task on these invariant images, their reference images, and MEI controls (Fig. 4a, top; for details see Section 3.2 and Supplementary Section B). Invariant images from the robust network (all generation stages) were significantly more recognizable by humans than their standard network counterparts (Fig. 4b, compare the green to the red boxes; p < 0.001, Supplementary Material Table  S2). Invariant images from the $L_{2}$-robust network were more recognizable when stretching was applied to earlier-layer representations, with pixel-space stretching yielding the highest discrimination accuracy. Interestingly, an opposite trend was observed for the standard network: invariant images from a late layer were more recognizable than those from a middle layer or from pixel space (p < 0.001 for each of the above comparisons: see Supplementary Material Table S2). These findings provide a more granular understanding of the perceptual misalignment between humans and CNNs, extending beyond what classic feature visualization tools can reveal with a direct comparison of MEIs. In fact, human participants classified the MEIs obtained for the robust network (Fig. 4b, cyan box) with an accuracy comparable to that achieved on the natural images encoded by the corresponding readout units (Fig. 4b, blue box). By comparison, human classification of the MEIs obtained for the standard network was relatively poor (Fig. 4b, pink box). Therefore, considering only MEIs’ interpretability, the $L_{2}$ robustification process would seem very effective in increasing the alignment of image classification between ResNet50 and human observers. SnS reveals instead that this alignment gradually deteriorates when considering invariant images obtained by stretching the representation in progressively deeper layers, with the interpretability gap between the robust and standard networks progressively narrowing. Interestingly, this also shows how SnS can uncover trends that are complementary to those found by previous approaches aimed at assessing invariance, such as computation of metamers (Feather et al., 2023). Similarly, SnS provides a complementary tool to test model alignment with human perception with respect to previous approaches like controversial stimuli (Golan et al., 2020) and Maximum differentiation competitions (Wang and Simoncelli, 2008), which study alignment through model disagreement. The same applies to other approaches that have examined human interpretability of parametric image transformations at the pixel or patch level, such as Extreme Image Transforms (EITs) (Malik et al., 2023) and Minimal Recognizable Configurations (MIRCs) (Ullman et al., 2016). An interesting question would be to assess how strongly the interpretability of SnS-generated invariant images correlates with the interpretability of EITs across a pool of generator networks, spanning different architectures and training regimes.
@@
To demonstrate the flexibility of the SnS framework, we performed the same interpretability experiment on invariant images evolved for readout units of a vision transformer (ViT). Similar to CNNs, the nature and interpretability of ViT invariances depended on the hierarchical level at which the representation was stretched. Invariances generated from pixel space were substantially less interpretable than those from mid- and high-level representations, which were remarkably similar and more interpretable. This observation aligns with the view that ViTs learn less strictly hierarchical and more globally-integrated features than CNNs (Dosovitskiy et al. (2020); see Fig. S9 in Section G of the Supplementary Material). Taken together, the results of Fig. 4 and Fig. S9 indicate that the level of transferability of the invariances obtained for the readout units of a generator network depends on three key factors: 1) the architecture of the generator (i.e., CNNs vs. vision transformers); 2) its training procedure (robustified vs. standard, in the case of CNNs); and 3) the hierarchical stage at which the representation is stretched to produce the invariant images.
@@ 5 Discussion
Limitations and broader impacts. While SnS has proven effective in uncovering novel, meaningful invariances, it presents opportunities for refinement and broader application. One avenue is to use SnS for further, more extensive geometric characterizations of the invariance manifolds, expanding the work presented in Supplementary Material, Section H. Another interesting direction is to use alternative generative models, such as large-scale GANs and diffusion models, to search the image space for invariant images. However, these models impose very strong priors on the structure of the synthesized images, which risks evolving exceedingly complex photorealistic details that are unlikely to be encoded by units in low- and mid-level processing stages (Wang and Ponce, 2024). Yet another possibility is to apply SnS to characterize invariance in other modalities, such as audition, using a generator capable of producing audio waveforms or spectrograms from a low-dimensional latent code - e.g., SpecGAN, WaveGAN (Donahue et al., 2018). Returning to the visual domain, our work has demonstrated SnS’ generative capabilities using eight different neural models (Fig. 4 and S9), belonging to three different architectural families, with parameter counts spanning roughly an order of magnitude (from 11.7M for ResNet-18 to 86.6M for ViT). This provides a strong test of our framework’s generalizability, although future work could test SnS with even larger models (e.g., DINO (Caron et al., 2021), ViT22-B (Dehghani et al., 2023)). Finally, by pinpointing where the perceptual misalignement between robustified networks and humans is strongest (i.e., in higher-order invariances), SnS could guide the development of models that achieve more human-aligned invariance. For instance, one could design a visual diet where “good” (i.e., human-like) and “bad” (i.e., nonhuman-like) invariant images are explicitly used in the training of the network.
@@ A.2 Experimental Hyperparameters
The generative model $\psi$ is instantiated as a pretrained deep neural network variant, specifically the fc7 configuration from Dosovitskiy and Brox (2016).
@@ Appendix E SnS is effective in targeting units in hidden layers
Appendix F SnS is robust to subsampling of the representation space
@@
Appendix G Hierarchical invariances in Vision Transformers
@@
Appendix H SnS can be used to study invariance manifold geometry
Appendix I SnS invariances and metamers are structurally and functionally distinct
@@
Appendix J SnS discovers layer specific hierarchical invariances for $L_{\infty}$-robust models

### v3

@@ 3.1 The Stretch-and-Squeeze algorithm (SnS)
For the characterization of invariance, we set this representation space at three distinct hierarchical levels within ResNet50: (i) $\kappa=0$, i.e., the input pixel space (denoted low_level), (ii) a mid-level convolutional layer (1st convolutional stage in layer 3, mid_level), and (iii) a deep convolutional layer (7th convolutional stage of layer 4, high_level). Our primary target units $u_{\mathrm{targ}}^{\ell}$ were chosen in the final readout layer (fully connected), as these units are selective for specific object categories, allowing us to study the invariance of the whole network to transformations of the image classes those neurons are tuned to. To demonstrate SnS’s generalizability beyond category-selective units, we also performed optimizations targeting units within mid_level and high_level, which are potentially more analogous to biological neurons in the visual system. In these cases, the input-side distance was computed in pixel space (i.e. $k=0$). For adversarial attack generation, we restricted our analysis to the configuration targeting readout layer units and defining input distance in pixel space. SnS optimizations were conducted on the units of two specific instances of ResNet50, our subject network: the standard ImageNet-pretrained model available in PyTorch (Paszke et al., 2019), and a robust counterpart generated via adversarial training with an $L_{2}$ perturbation norm constraint of $\epsilon=3$ (Engstrom et al., 2019a). SnS invariances were studied also in ResNet18 and VGG16_bn (see Section 3.2) and Vision Transformers (ViT, see Supplementary Material, Section I).
@@ 4.2 SnS reveals layer-specific invariant manifolds
Yet another way to characterize the invariance manifolds is to estimate their intrinsic dimension (ID). We did so for an example readout unit of $L_{2}$-robust ResNet50, using both PCA and 2NN-ID, a state-of-the-art nonlinear estimator (Facco et al., 2017) (see Section F in the Supplementary Material). 2NN-ID yielded ID estimates one order of magnitude lower than PCA, indicating that the geometry of the three manifolds was highly nonlinear (Fig. S8). Both methods, however, revealed a consistent ID ranking across stretching layers: lowest for the pixel space, highest for mid layers, and lower again for deep layers - a pattern mirroring previously reported trends in the ID of image representations across deep CNNs and visual cortex (Ansuini et al., 2019; Muratore et al., 2022a)
@@
Hierarchical invariances are robust to representation subsampling. To assess SnS’s applicability to neuroscience, where only a small neuronal subpopulation is typically recorded from a visual area, we generated the invariant images using heavily subsampled hidden layer representations. Both qualitative inspection and representational distance analysis revealed that the resulting invariances closely resembled those obtained using the full layer representation, supporting SnS’s potential for neuroscience experiments (see Supplementary Material, Section G for further details).
@@ 4.3 Invariant images for robust and standard networks are perceptually different
We next evaluated how recognizable the invariant images generated by SnS were by humans and other CNNs. In our tests, invariant images for both kinds of networks (i.e., standard and $L_{2}$-robust) were generated by stretching the low- (pixel), mid- and high-level representations of the same set of natural images sampled from 12 different Imagenet categories (Fig. 4a). Human subjects performed a 12-AFC classification task on these invariant images, their reference images, and MEI controls (Fig. 4a, top; for details see Section 3.2 and Supplementary Section B). Invariant images from the robust network (all generation stages) were significantly more recognizable by humans than their standard network counterparts (Fig. 4b, compare the green to the red boxes; p < 0.001, Supplementary Material Table  S2). Invariant images from the $L_{2}$-robust network were more recognizable when stretching was applied to earlier-layer representations, with pixel-space stretching yielding the highest discrimination accuracy. Interestingly, an opposite trend was observed for the standard network: invariant images from a late layer were more recognizable than those from a middle layer or from pixel space (p < 0.001 for each of the above comparisons: see Supplementary Material Table S2). These findings provide a more granular understanding of the perceptual misalignment between humans and CNNs, extending beyond what classic feature visualization tools can reveal with a direct comparison of MEIs. In fact, human participants classified the MEIs obtained for the robust network (Fig. 4b, cyan box) with an accuracy comparable to that achieved on the natural images encoded by the corresponding readout units (Fig. 4b, blue box). By comparison, human classification of the MEIs obtained for the standard network was relatively poor (Fig. 4b, pink box). Therefore, considering only MEIs’ interpretability, the $L_{2}$ robustification process would seem very effective in increasing the alignment of image classification between ResNet50 and human observers. SnS reveals instead that this alignment gradually deteriorates when considering invariant images obtained by stretching the representation in progressively deeper layers, with the interpretability gap between the robust and standard networks progressively narrowing. Interestingly, this also shows how SnS can uncover trends that are complementary to those found by previous approaches aimed at assessing invariance, such as computation of metamers (Feather et al., 2023). Similarly, SnS provides a complementary tool to test model alignment with human perception with respect to previous approaches which study alignment through model disagreement (Wang and Simoncelli, 2008; Golan et al., 2020), as well as to other approaches that have examined human interpretability of parametric image transformations at the pixel or patch level (Ullman et al., 2016; Malik et al., 2023).
@@
To demonstrate the flexibility of the SnS framework, we performed the same interpretability experiment on invariant images evolved for readout units of a vision transformer (ViT). Similar to CNNs, the nature and interpretability of ViT invariances depended on the hierarchical level at which the representation was stretched. Invariances generated from pixel space were substantially less interpretable than those from mid- and high-level representations, which were remarkably similar and more interpretable. This observation aligns with the view that ViTs learn less strictly hierarchical and more globally-integrated features than CNNs (Dosovitskiy et al. (2020); see Fig. S11 in Section I of the Supplementary Material). Taken together, the results of Fig. 4 and Fig. S11 indicate that the level of transferability of the invariances obtained for the readout units of a generator network depends on three key factors: 1) the architecture of the generator (i.e., CNNs vs. vision transformers); 2) its training procedure (robustified vs. standard, in the case of CNNs); and 3) the hierarchical stage at which the representation is stretched to produce the invariant images.
@@ 5 Discussion
Limitations and broader impacts. While SnS has proven effective in uncovering novel, meaningful invariances, it presents opportunities for refinement and broader application. One avenue is to use SnS for further, more extensive geometric characterizations of the invariance manifolds, expanding the work presented in Supplementary Material, Section F. Another interesting direction is to apply SnS to characterize invariance in other modalities, such as audition, using a generator capable of producing audio waveforms or spectrograms from a low-dimensional latent code - e.g., SpecGAN, WaveGAN (Donahue et al., 2018). Returning to the visual domain, our work has demonstrated SnS’ generative capabilities using eight different neural models (Fig. 4 and S11), belonging to three different architectural families, with parameter counts spanning roughly an order of magnitude (from 11.7M for ResNet-18 to 86.6M for ViT). This provides a strong test of our framework’s generalizability, although future work could test SnS with even larger models (e.g., DINO (Caron et al., 2021), ViT22-B (Dehghani et al., 2023)). Finally, by pinpointing where the perceptual misalignement between robustified networks and humans is strongest (i.e., in higher-order invariances), SnS could guide the development of models that achieve more human-aligned invariance. For instance, one could design a visual diet where “good” (i.e., human-like) and “bad” (i.e., nonhuman-like) invariant images are explicitly used in the training of the network.
@@ 7 Ethics statement
8 Acknowledgements
We would like to thank Sebastiano Quintavalle for his help in implementing the code, and Eugenio Piasini for his suggestions on the Pareto front optimization. This work was funded by the European Union – NextGenerationEU – NRRM4C2-I.1.1, in the framework of the PRIN Project no. 2022WX3FM5, CUP:G53D23003220006 (DZ), by NIGMS Award T32GM144273 (MT), by NIH grant R01EY026025 (GK), and by the NSF Center for Brains, Minds, and Machines NSF CCF-1231216 (GK). The content of this paper is solely the responsibility of the authors and does not necessarily represent the official views of any of the above organizations.
@@ A.2 Experimental Hyperparameters
The generative model $\psi$ is instantiated as a pretrained deep neural network variant, specifically the fc7 configuration from Dosovitskiy and Brox (2016). The choice of this model instead of alternative models, such as as large-scale GANs and diffusion models, was motivated by the fact that the latter impose very strong priors on the structure of the synthesized images, which risks evolving exceedingly complex photorealistic details that are unlikely to be encoded by units in low- and mid-level processing stages (Wang and Ponce, 2024).
@@ Appendix E SnS is effective in targeting units in hidden layers
Appendix F SnS can be used to study invariance manifold geometry
Appendix G SnS is robust to subsampling of the representation space
@@
Appendix H SnS discovers layer specific hierarchical invariances for $L_{\infty}$-robust models
Appendix I Hierarchical invariances in Vision Transformers
@@
Appendix J SnS invariances and metamers are structurally and functionally distinct